a number when multiplied by 7/18 instead of 7/8 and got the result 770 less than the actual result , find the original number?

Grade:

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

let no is x

x*7/18 = x*7/8 - 770

solving x=11*8*18=1584

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Suppose the number be x

x*7/18 = x*7/8 - 770

hence,

x=11*8*18=1584

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi